# DM-MPPI: Datamodel for Efficient and Safe Model Path Integral Control

### v1

@@ Abstract
We extend the Datamodels framework from machine learning to Model Predictive Path Integral (MPPI) control. While datamodels estimate sample influence via regression on fixed datasets, our approach learns to predict influence from sample cost features—enabling real-time estimation for new samples without online regression. The predictor is trained offline on influence coefficients computed via the datamodel framework across diverse MPPI instances, then deployed for efficient sample pruning and adaptive constraint handling. A single learned model addresses both efficiency and safety: low-influence samples are pruned to reduce computation, while monitoring constraint-violation influence enables adaptive penalty tuning. Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining performance and improving constraint satisfaction.
@@ I Introduction
Model Predictive Path Integral (MPPI) control has emerged as a powerful framework for sampling-based optimal control in nonlinear systems [1, 2]. By generating trajectory samples and computing importance-weighted averages, MPPI avoids linearization while achieving real-time performance through parallelization [3]. However, two fundamental challenges remain. First, sample efficiency: many samples may contribute little to the final control, wasting computation. Second, constraint handling: soft constraint penalties are the standard approach [18], but understanding how constraint-violating samples influence the control output remains opaque.
The Datamodels framework from machine learning [8] offers a principled approach to understanding how individual data points affect model outputs. We adapt this framework to MPPI by learning to predict trajectory cost from sample inclusion. The key insight is simple: by predicting total cost, which includes constraint violation penalties, a single datamodel captures both task performance and constraint-relevant influence. The relative importance of constraints can be adjusted through the penalty coefficient without requiring separate models.
I-A Related Work
MPPI and Path Integral Control. Path integral methods derive optimal control laws through exponential weighting of sampled trajectories [4]. Williams et al. [1, 2] developed MPPI for real-time control with GPU parallelization, demonstrating aggressive autonomous driving [3]. Recent variants address robustness [5], smoothness [6], and sample efficiency through unscented guidance [7]. Our work complements these advances by providing interpretable influence analysis.
Data Attribution in Machine Learning. Understanding how training data affects model predictions is a fundamental problem in ML. Influence functions [9] adapt classical robust statistics to deep learning, while Data Shapley [10] applies game-theoretic valuation. TracIn [11] provides efficient gradient-based attribution. The Datamodels framework [8] learns linear predictors of model outputs from training set composition, achieving high accuracy on image classification. TRAK [12] scales attribution to large models via random projections. We are the first to apply datamodels to sampling-based control.
Constraint Handling in MPPI. Standard MPPI handles constraints through soft penalties in the cost function, but this approach cannot guarantee constraint satisfaction and requires careful tuning of penalty weights. To address this limitation, several MPPI-specific methods have been developed. Risk-Aware MPPI (RA-MPPI) [13] replaces the expected cost with Conditional Value-at-Risk (CVaR), steering optimization toward the worst-case tail of the rollout distribution to improve safety in autonomous racing. Shield-MPPI [14] integrates Control Barrier Functions (CBFs) into MPPI through a dual-layer structure that combines cost penalization with a gradient-based local repair step, significantly reducing constraint violations while maintaining computational efficiency on CPUs. For hard constraint satisfaction, Constrained Covariance Steering MPPI (CCS-MPPI) [15] enforces state and control constraints through a tube-based formulation, though it operates probabilistically. Gandhi et al. [22] develop safe importance sampling that biases trajectory generation toward constraint-satisfying regions while preserving theoretical guarantees. Our datamodel-based approach complements these methods by diagnosing how constraint-violating samples influence the control output, enabling principled tuning of penalty parameters rather than enforcing constraints directly.
Learning for Control. Deep learning increasingly intersects with control theory [23, 24]. Data-driven methods can directly synthesize controllers from data [25]. Neural certificates [26] learn Lyapunov and barrier functions alongside policies. Interpretability of learned policies is an emerging concern [27]. Our work contributes to this intersection by bringing data attribution tools to sampling-based control.
I-B Contributions
This paper extends the Datamodels framework to MPPI by learning to predict sample influence from cost features rather than computing influence for fixed datasets. Our key contribution is an offline-online architecture: ground-truth influence coefficients are computed via datamodel regression during training, then predicted via a neural network at runtime—enabling real-time estimation without online regression. A single learned predictor addresses both sample efficiency through pruning and constraint handling through adaptive penalty tuning.
The remainder of this paper is organized as follows. Section II provides background on MPPI control and the Datamodels framework. Section III presents the proposed MPPI-Datamodel framework, detailing the offline training procedure for learning the influence predictor and the online inference phase for sample pruning and adaptive constraint handling. Section IV evaluates the framework on a path tracking task with obstacle avoidance, demonstrating sample efficiency gains and improved safety margins. Section V discusses limitations and directions for future work, and Section VI concludes the paper.
@@ II-A Model Predictive Path Integral Control
Consider a nonlinear stochastic system with control-affine dynamics
$dx=f(x)dt+G(x)(u+\epsilon)dt$ (1)
where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{m}$ is the control input, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the drift dynamics, $G:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ is the control matrix, and $\epsilon\sim\mathcal{N}(0,\Sigma)$ represents control noise.
@@
where the goal cost captures task objectives $S_{\text{goal}}(\tau)=\phi(x_{T})+\sum_{t=0}^{T-1}q(x_{t})$, the control cost penalizes actuation $S_{\text{control}}(\tau)=\sum_{t=0}^{T-1}\frac{1}{2}u_{t}^{\top}Ru_{t}$, and the violation cost captures soft constraint penalties $S_{\text{violation}}(\tau)=\sum_{t=0}^{T-1}c(x_{t})$ with $\rho>0$ controlling constraint importance.
@@ II-B The Datamodels Framework
where $\mathbf{1}_{S^{\prime}}\in\{0,1\}^{n}$ is a binary vector indicating which data points are included in subset $S^{\prime}$. The coefficient $\theta_{j}$ quantifies the marginal influence of data point $z_{j}$ on the output: positive $\theta_{j}$ indicates that including $z_{j}$ increases the predicted output, while negative $\theta_{j}$ indicates a decreasing effect.
Training proceeds by constructing many random subsets $\{S^{\prime}_{i}\}_{i=1}^{M}$, each formed by including each data point independently with probability $\alpha\in(0,1)$. For each subset, the true model output $f(x;S^{\prime}_{i})$ is computed. The datamodel parameters are then learned via LASSO regression [12]:
@@
II-C Why Linear Models Suit MPPI
The weighted cost can be written as
$\bar{S}(\mathbf{b})=\frac{\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}}{\sum_{j}b_{j}e^{-S_{j}/\lambda}}=\frac{N(\mathbf{b})}{D(\mathbf{b})}$ (7)
where both numerator $N(\mathbf{b})$ and denominator $D(\mathbf{b})$ are linear in the inclusion vector $\mathbf{b}$, but their ratio is not. However, when subsets are formed by including each sample independently with probability $\alpha$, the denominator concentrates around its expectation for sufficiently large $\alpha K$:
$D(\mathbf{b})\approx\mathbb{E}[D]=\alpha\sum_{k}e^{-S_{k}/\lambda}.$ (8)
Substituting this approximation yields
$\bar{S}(\mathbf{b})\approx\frac{1}{\mathbb{E}[D]}\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}=\sum_{k}\theta_{k}b_{k}$ (9)
where $\theta_{k}\propto e^{-S_{k}/\lambda}S_{k}$. Thus, when subset size is large enough for the denominator to concentrate, the weighted cost becomes approximately linear in $\mathbf{b}$. This approximation degrades when $\alpha K$ is small or when a few samples dominate the weights (low $\lambda$, high cost variance), causing $D(\mathbf{b})$ to fluctuate significantly across subsets. Nonetheless, even when these conditions are not fully met, the linear datamodel can still serve as a practical approximation worth investigating.
@@ III Method: MPPI-Datamodel Framework
[…]
Figure 1: MPPI-Datamodel workflow. Offline phase (green): collect MPPI instances, fit datamodels via LASSO regression, and train influence predictor $h_{\phi}$. Online phase: predict influence using $h_{\phi^{*}}$ with cost features and instance statistics, prune low-influence samples for efficiency, and adapt penalty $\rho$ based on violation influence ratio for safety.
@@ III-A Problem Formulation
where $\rho>0$ is the soft constraint penalty coefficient. For a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$, the subset-restricted weighted cost is
$\bar{C}(\mathcal{T}^{\prime})=\sum_{k:\tau_{k}\in\mathcal{T}^{\prime}}w_{k}(\mathcal{T}^{\prime})\,C_{k}$ (11)
with renormalized importance weights
@@
To capture the relationship between sample inclusion and weighted cost, we define the binary inclusion vector $\mathbf{b}\in\{0,1\}^{K}$ where $b_{k}=1$ if trajectory $\tau_{k}$ is included. The MPPI-Datamodel predicts weighted cost from sample membership via the linear model
$g_{\theta}(\mathbf{b})=\theta^{\top}\mathbf{b}+\theta_{0}\approx\bar{C}(\mathcal{T}^{\prime})$ (13)
where the coefficient $\theta_{k}$ represents the marginal influence of sample $\tau_{k}$ on the weighted cost. Since the MPPI control update $u^{*}=\sum_{k}w_{k}u_{k}$ uses the same importance weights $w_{k}$, samples that significantly affect the weighted cost also significantly affect the control output. This justifies using influence on weighted cost as a proxy for influence on the control decision.
@@ III-B Offline Training Phase
Step 2: Datamodel Fitting. For each instance $i$, we learn the influence coefficients by regression on random subsets. We generate $M$ random subsets by including each trajectory independently with probability $\alpha\in(0,1)$. For each subset $j$ with inclusion mask $\mathbf{b}_{j}\in\{0,1\}^{K}$, we compute the subset-restricted weighted cost
$\bar{C}_{j}^{(i)}=\frac{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}}$ (14)
where $w_{k}^{(i)}=\exp(-C_{k}^{(i)}/\lambda)$ are the unnormalized importance weights. The influence coefficients are obtained by solving the LASSO regression problem
$\theta^{(i)*}=\operatorname*{arg\,min}_{\theta,\theta_{0}}\frac{1}{M}\sum_{j=1}^{M}\left(\theta^{\top}\mathbf{b}_{j}+\theta_{0}-\bar{C}_{j}^{(i)}\right)^{2}+\mu\|\theta\|_{1}$ (15)
where $\theta_{0}$ is an intercept term excluded from regularization and the $\ell_{1}$ penalty encourages sparsity.
@@
This independence decorrelates the contributions of different samples, allowing the regression to disentangle individual effects from aggregate subset effects. With sufficient subsets $M$, the LASSO solution $\theta_{k}^{(i)*}$ accurately estimates the marginal influence of each sample $k$.
@@
To capture this context dependence, we augment the input features with summary statistics:
$\hat{\theta}_{k}=h_{\phi}\left(C_{k},C_{\text{viol},k},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)$ (17)
where $\bar{C}^{(i)}=\frac{1}{K}\sum_{k=1}^{K}C_{k}^{(i)}$ is the mean cost and $\sigma_{C}^{(i)}=\text{std}(\{C_{k}^{(i)}\}_{k=1}^{K})$ is the standard deviation. These statistics provide context about where sample $k$ sits relative to other samples in the instance.
The predictor parameters are learned by minimizing prediction error across all collected instances:
$\phi^{*}=\operatorname*{arg\,min}_{\phi}\sum_{i=1}^{N}\sum_{k=1}^{K}\left(h_{\phi}\left(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)-\theta_{k}^{(i)*}\right)^{2}$ (18)
@@
10:   Compute weighted cost: $\bar{C}_{j}^{(i)}\leftarrow\frac{\sum_{k}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k}b_{j,k}\,w_{k}^{(i)}}$
11:  end for
12:  Solve LASSO (15) to obtain $\theta^{(i)*}\in\mathbb{R}^{K}$
13:  for $k=1$ to $K$ do
14:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)},\theta_{k}^{(i)*})\}$
15:  end for
16:end for
17: Train $h_{\phi^{*}}$ on $\mathcal{D}$ by minimizing (18)
18:return $h_{\phi^{*}}$

### v2

# DM-MPPI: Datamodel-Based Sample Pruning for Model Predictive Path Integral Control

@@ Abstract
Sampling-based controllers like Model Predictive Path Integral (MPPI) generate hundreds of trajectory rollouts every control cycle, but many of these rollouts barely affect—or even hurt—the final control action. Figuring out which samples actually matter is, at its core, a data attribution problem: each rollout acts like a training example that pushes the controller’s output one way or another, depending on whether it is included. We make this connection concrete by adapting the Datamodels framework, which learns how much each sample matters through regression over random subsets, to MPPI. During an offline phase, we compute influence scores across random subsets of rollouts and train a small neural network to predict influence from cost features—allowing real-time pruning without any online regression. An adaptive constraint penalty also adjusts obstacle avoidance strength based on how often sampled trajectories violate constraints. We validate the method on a kinematic bicycle model driving through two obstacle-filled tracks.
@@ I Introduction
Model Predictive Path Integral (MPPI) control has become a popular sampling-based approach for nonlinear systems [17, 16], achieving real-time performance by rolling out many trajectories in parallel and blending them through importance-weighted averaging [18]. However, two key challenges remain: a large fraction of the sampled trajectories has little effect on the final control action, leading to wasted computation, and soft constraint penalties [23] make it unclear how constraint-violating samples shape the output. A natural question is whether we can identify which samples actually matter and focus the computation on those. We argue that this is naturally a data attribution problem — each rollout is like a training example that, when included or left out, shifts the control update in a different direction. The Datamodels framework [7] offers a clean solution by learning per-example influence coefficients through subset regression. We bring this idea into MPPI by moving the expensive regression to an offline stage and training a small neural network to predict influence at deployment time.
Path integral methods recover optimal control laws by exponentially weighting sampled trajectories [14]. Williams et al. [17, 16] introduced MPPI for real-time GPU-parallel control and showed its effectiveness in aggressive autonomous driving [18]. Follow-up work has since improved robustness [5], smoothness [8], and sample efficiency [11]. On the constraint side, researchers have explored CVaR-based risk awareness [22], Control Barrier Functions [21], tube-based formulations [20], and safe importance sampling [4]. Deep learning intersects with control through data-driven synthesis [3, 10, 19] and neural certificates [2], with interpretability drawing growing attention [15]. In the machine learning literature, data attribution has been studied through influence functions [9], Data Shapley [6], TracIn [13], the Datamodels framework [7], and TRAK [12]. To our knowledge, no prior work has applied data attribution to sampling-based control.
Our contributions are threefold. First, we recast MPPI sample selection as a data attribution problem and adapt the Datamodels framework through an offline–online setup: influence coefficients are fitted via subset regression during training and predicted by a neural network at runtime. Second, we compare nine methods on a kinematic bicycle model across two track geometries with obstacles. Keeping only $20\%$ of samples through informed pruning substantially lowers tracking error compared to unpruned MPPI, while random pruning is worse than even the smallest sample budget. Adding more samples shows diminishing returns, but pruning down to a few well-chosen samples consistently beats the full unpruned budget. Third, we show that the learned influence predictor with adaptive mechanisms (DM-adapted) achieves the fewest boundary violations across both tracks, significantly outperforming heuristic pruning methods that tend to keep trajectories running tight along obstacles—whereas influence-aware selection naturally avoids putting too much weight on samples near the constraint boundary.
The rest of this paper is organized as follows. Section II covers background on MPPI and the Datamodels framework. Section III presents the proposed framework. Section IV evaluates against baselines on two track geometries. Section V concludes.
@@ II-A Model Predictive Path Integral Control
Consider a discrete-time nonlinear system with a control-affine structure.
$x_{t+1}=F(x_{t},u_{t}+\epsilon_{t})$ (1)
where $x_{t}\in\mathbb{R}^{n}$ is the state, $u_{t}\in\mathbb{R}^{m}$ is the control input, $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is the dynamics map, and $\epsilon_{t}\sim\mathcal{N}(0,\Sigma)$ represents zero-mean Gaussian control noise with covariance $\Sigma\in\mathbb{R}^{m\times m}$.
@@
where the goal cost captures task objectives $S_{\text{goal}}(\tau)=\phi(x_{T})+\sum_{t=0}^{T-1}q(x_{t})$, the control cost follows the standard MPPI formulation $S_{\text{control}}(\tau_{k})=\lambda\sum_{t=0}^{T-1}\bar{u}_{t}^{\top}R\,\epsilon_{t,k}$ (a cross-term between the nominal control and the perturbation [16]), and the violation cost captures soft constraint penalties $S_{\text{violation}}(\tau)=\sum_{t=0}^{T-1}c(x_{t})$ with $\rho>0$ controlling constraint importance.
@@ II-B The Datamodels Framework
where $\mathbf{1}_{S^{\prime}}\in\{0,1\}^{n}$ is a binary vector indicating which data points are included in the subset $S^{\prime}$. The coefficient $\theta_{j}$ quantifies the marginal influence of data point $z_{j}$ on the output: a positive $\theta_{j}$ indicates that including $z_{j}$ increases the predicted output, while a negative $\theta_{j}$ indicates a decreasing effect.
Training proceeds by constructing many random subsets $\{S^{\prime}_{i}\}_{i=1}^{M}$, each formed by including each data point independently with a probability $\alpha\in(0,1)$. For each subset, the true model output $f(x;S^{\prime}_{i})$ is computed. The datamodel parameters are then learned via LASSO regression [12]:
@@
II-C Motivation for Linear Influence Models in MPPI
The MPPI control update (3) is a weighted sum of perturbations, where inclusion or exclusion of a sample changes both the numerator and the normalizing denominator of the weights. The control update is therefore not linear in the inclusion vector $\mathbf{b}\in\{0,1\}^{K}$, and there is no formal guarantee that the LASSO regression will recover accurate influence coefficients. Nonetheless, empirically we find that the linear datamodel achieves $R^{2}\approx 0.6$ when predicting the first-step control update from sample inclusion (Section IV). We treat this as a practical heuristic: the moderate $R^{2}$ indicates that a linear model captures some structure in sample influence, and the MLP predictor trained on these coefficients inherits both the signal and the approximation error. Whether the resulting pruning criterion is better than simple cost-based heuristics is an empirical question that we investigate in Section IV.
@@ III Method: MPPI-Datamodel Framework
Figure 1: DM-MPPI framework.
@@ III-A Problem Formulation
where $\rho>0$ is the soft constraint penalty coefficient. For a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$, the subset-restricted control update is computed by renormalizing importance weights over $\mathcal{T}^{\prime}$:
@@
$\delta u_{0}(\mathcal{T}^{\prime})=\sum_{k:\tau_{k}\in\mathcal{T}^{\prime}}w_{k}(\mathcal{T}^{\prime})\,\epsilon_{0,k}$ (9)
where $\epsilon_{0,k}$ is the first-step control perturbation of trajectory $\tau_{k}$.
To capture the relationship between sample inclusion and the control update, we define the binary inclusion vector $\mathbf{b}\in\{0,1\}^{K}$ where $b_{k}=1$ if trajectory $\tau_{k}$ is included. The MPPI-Datamodel predicts the control update from sample membership via the linear model
$g_{\theta}(\mathbf{b})=\theta^{\top}\mathbf{b}+\theta_{0}\approx\delta u_{0}(\mathcal{T}^{\prime})$ (10)
where we fit separate regressions for each control dimension and average the absolute coefficients: $|\theta_{k}|=\frac{1}{m}\sum_{d=1}^{m}|\theta_{k}^{(d)}|$. The resulting $|\theta_{k}|$ quantifies the marginal influence of sample $\tau_{k}$ on the immediate control action. Samples with large $|\theta_{k}|$ disproportionately affect the first control applied to the system.
@@ III-B Offline Training Phase
Step 2: Datamodel Fitting. For each instance $i$, we learn the influence coefficients by regression on random subsets. We generate $M$ random subsets by including each trajectory independently with probability $\alpha\in(0,1)$. For each subset $j$ with an inclusion mask $\mathbf{b}_{j}\in\{0,1\}^{K}$, we compute the subset-restricted control update $\delta u_{0,j}^{(i)}\in\mathbb{R}^{m}$ by renormalizing the MPPI weights over the included trajectories and taking the weighted sum of first-step perturbations. The influence coefficients for each control dimension $d$ are obtained by solving
$\theta^{(i,d)*}=\operatorname*{arg\,min}_{\theta,\theta_{0}}\frac{1}{M}\sum_{j=1}^{M}\left(\theta^{\top}\mathbf{b}_{j}+\theta_{0}-[\delta u_{0,j}^{(i)}]_{d}\right)^{2}+\mu\|\theta\|_{1}$ (11)
where $\theta_{0}$ is an intercept term excluded from regularization, and the $\ell_{1}$ penalty encourages sparsity. We then compute a single influence score per sample by averaging absolute coefficients across dimensions: $|\theta_{k}^{(i)*}|=\frac{1}{m}\sum_{d=1}^{m}|\theta_{k}^{(i,d)*}|$.
@@
This independence decorrelates the contributions of different samples, helping the regression disentangle individual effects from aggregate subset effects. The LASSO solution $\theta_{k}^{(i)*}$ provides an approximate influence score for each sample $k$; the quality of this approximation depends on $M$ and on the degree of nonlinearity in the control update mapping (Section II-C).
@@
To capture this context dependence, we construct an 8-dimensional feature vector:
$\hat{\theta}_{k}=h_{\phi}\left(\mathbf{f}_{k}^{(i)}\right)$ (13)
where $\mathbf{f}_{k}^{(i)}=[C_{k},\;C_{\text{viol},k},\;z_{k},\;\tilde{C}_{k},\;s_{k},\;\mathbf{1}_{C_{\text{viol},k}>0},\;\bar{C}^{(i)},\;\sigma_{C}^{(i)}]^{\top}$. Here $z_{k}=(C_{k}-\bar{C})/\sigma_{C}$ is the $z$-score, $\tilde{C}_{k}=(C_{k}-C_{\min})/\sigma_{C}$ is the normalized distance from the best sample, $s_{k}=\exp(-(C_{k}-C_{\min})/\sigma_{C})$ is a soft exponential rank, $\bar{C}^{(i)}=\frac{1}{K}\sum_{k}C_{k}^{(i)}$ is the mean cost, and $\sigma_{C}^{(i)}=\text{std}(\{C_{k}^{(i)}\})$ is the cost standard deviation. The per-sample features ($z_{k},\tilde{C}_{k},s_{k}$) encode where sample $k$ sits in the cost distribution, while the instance statistics ($\bar{C},\sigma_{C}$) provide context.
The predictor parameters are learned by minimizing the prediction error across all collected instances:
$\phi^{*}=\operatorname*{arg\,min}_{\phi}\sum_{i=1}^{N}\sum_{k=1}^{K}\left(h_{\phi}\left(\mathbf{f}_{k}^{(i)}\right)-|\theta_{k}^{(i)*}|\right)^{2}$ (14)
@@
10:   Compute subset control update: $\delta u_{0,j}^{(i)}\leftarrow\sum_{k:b_{j,k}=1}\tilde{w}_{k}^{(i,j)}\,\epsilon_{0,k}^{(i)}$
11:    with $\tilde{w}_{k}^{(i,j)}=\exp(-C_{k}^{(i)}/\lambda)/\sum_{l:b_{j,l}=1}\exp(-C_{l}^{(i)}/\lambda)$
12:  end for
13:  Solve LASSO (11) per control dimension, average $|\theta_{k}^{(i)*}|=\frac{1}{m}\sum_{d}|\theta_{k}^{(i,d)*}|$
14:  for $k=1$ to $K$ do
15:   Build feature vector $\mathbf{f}_{k}^{(i)}$ per (13)
16:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{f}_{k}^{(i)},|\theta_{k}^{(i)*}|)\}$
17:  end for
18:end for
19: Train $h_{\phi^{*}}$ on $\mathcal{D}$ by minimizing (14)
20:return $h_{\phi^{*}}$